If and object of 80kg falls from a height of 90m what is the velocity the object will fall with?


1
Expert's answer
2021-11-24T19:10:54-0500

We can find the velocity of the object with which it hits the ground from the law of conservation of energy:


"KE=PE,""\dfrac{1}{2}mv^2=mgh,""v=\sqrt{2gh}=\sqrt{2\times9.8\ \dfrac{m}{s^2}\times90\ m}=42\ \dfrac{m}{s}."